The length is x-4 and the width is x+6. Express the perimeter of the rectangle as a polynomial in simplest form, and express the area of the rectangle as a polynomial in simplest form
Answer by richard1234(7193)

Perimeter = 2L + 2W = 2(x-4) + 2(x+6) = 4x + 4

Area = LW = (x-4)(x+6) = x^2 + 6x - 4x - 24 = x^2 + 2x - 24